The QEs were measured at 293K, 239K, 232K, and 293K again. The cooling was provided by the PD TEC. At each temperature, the incident power was changed from 30uW to 1mW to see the dependence of the QE on the incident power to check the possible saturation.
The QE was 79~81% (the window T=96.6% was already compensated). I'm not 100% sure this 1% variation in the plateau is real or due to insufficient calibration of the REF PD.
The REF PD was calibrated at 1mW at 100mA injection current to the laser.
No obvious saturation was observed.
We can cool the PD with LN2 and we should make a careful alignment of the beam at each temperature. |